UARS Weekly Status Report

01 December 2000




********************************************************************

                UARS WEEKLY STATUS REPORT - 01 DEC 2000

********************************************************************

***********
SOLSTICE
***********

SOLSTICE is in good health, collecting solar data on all available orbital
tracking periods and stellar calibration observations whenever possible.
Our solar data is unaffected by the larger than normal pointing errors
since they are within our calibrations. Today (00336) we began day only
operations in support of the yaw manuever and will resume normal
operations  in two days. Additionally, we continue to gather enough data to
meet all the SOLSTICE science objectives during this period of reduced tape
recorder operations.

In support of the OA_SSPP_ATT patch testing we are running our modified
processing code under RSS. We are processing the last year's worth of data
in this manner and expect to validate this patch in the next week.

***********
SUSIM
***********

SUSIM performed daily solar scans of the sun during periods
of predicted telemetry.

We are preparing to begin the reprocessing the SUSIM UV spectral
irradiances to data version V21 in the next few weeks.

***********
ACRIM
***********

Nothing received.

***********
PEM
***********

Nothing received.

***********
HALOE 
***********

Nothing received.

***********
MLS
***********

Nothing received.

***********
HRDI
***********

Nothing received.

***********
WINDII
***********

  In order to conserve our human resources the WINDII team has decided
to cease WINDII operations for a few months in order to concentrate
on the analysis of recent data. Since we are now running without
active thermal control the WINDII operating temperature is about
8 C rather than the 20 C it was designed and calibrated for. Thus
we have had to recharacterize the instrument at this temperature, 
in flight, and that has taken more effort than expected. 
  We plan to begin operations again after the TIMED launch, because
the opportunity for simultaneous measurements at different local
times with TIMED should be very rewarding. 
  WINDII has just completed a (major) infrequent calibration, and
will thus now go into standby until further notice.  

***********
MPG
***********

Daily processing of instrument Daily Activity Plans continues 
with no problems.  The MPG continues to coordinate with the 
Flight Operations Team and to assist the Project Scientist 
with the scheduling of instrument operations.  The latest
published instrument operations schedule is complete to 3 January
2001. 

***********
CDHF
***********

Nothing received.
                  
***********
FOT
***********

                          REPORTING PERIOD

This report covers 18 November 2000 (Orbit 50221, GMT Day 323) through 24 
November 2000 (Orbit 50325, GMT Day 329).
                       
                        SPACECRAFT OPERATION

The observatory is now in Reverse Flight (Northern Hemisphere viewing) and is 
performing nominally under two-battery operations.  The instrument operational 
changes were:


                  INSTRUMENT OPERATIONAL CHANGES


     Instrument     Time       Orbit     Comment
 PEM          324/2325     50250     Off
 HALOE        325/0212     50252     On


The beta angle ranged between the angles of 56.5 degrees to 32.8 degrees for 
this report period.  The beta angle is now decreasing towards a minimum angle 
of 2 degrees on 3 December 2000 (DOY 337).

Spacecraft battery 1, 2, and 3 performance monitoring continues.  Battery 1 
remains OFF the charge relay.  Battery 2 and Battery 3 are currently 
maintaining greater than 23.8 V end-of-night (EON) load bus voltage.  Battery 
temperatures are stable with a temperature delta between Battery 2 and Battery 
3 of 1.3 and 1.2 degrees C respectively.  Additional battery performance data 
for each battery on 17 November 2000 and 24 November 2000 are provided in 
Appendix D.

The clock error ranged from 11.7 to 10.3 msec during this report period.  
Clock rate adjustments are listed in Appendix C.

All instrument and subsystem engineers were advised on the initial excursion of 
ALERTS or Out-Of-Limits conditions detected by the FOT for this report.  A 
summary is listed below and the details of ALERT and Out-Of-Limits occurrences 
are listed in Appendix B.

     OUT-OF-LIMITS  =  01
     ALERTS         =  00

The following real-time command sheets were executed for the listed operational 
element(s) during this report period (see Appendix A).


     FOT           =  23
     HALOE         =  03
     PEM           =  01

UARS Anomaly Reports initiated during this report period are listed in Appendix 
E and summarized below.

     FOT          =   01

One (1) Anomaly Report remained OPEN at the end of this report period (see 
Appendix E).

TRANSPONDER FREQUENCIES:
 Transponder A:  00/221  -  2287.496937
 Transponder B:  00/305  - 2287.499087
 
 
                       GROUND SYSTEM OPERATION
 
 One (1) operational support problem occurred this period.  Anomaly Reports are 
included in Appendix E.
 
The data loss calculations supplied by Data Capture Facility (DCF) from mission 
start through UTC Day 00/328 show the observatory data loss to be 299 hours, 58 
minutes, 58.600 seconds. This shows an overall decrease of 1 hour, 7 minutes 
and 53.016 seconds since the last report period due to an apparent data 
recovery from the previous week.  This is a 0.39 percent data loss, which 
equals a 99.61 percent data capture for the mission.
 
 The increase of 3 minutes and 53.736 seconds occurred over three (3) days:
 
 1 minute, 37.504 seconds on Day 322 due to a data hit
 0 minutes, 05.168 seconds on Day 323 due to a data hit
 2 minutes, 11.064 seconds on Day 325 due to a data hit


                    OTHER SIGNIFICANT EVENTS

                    UARS SATELLITE OPERATIONS

Testing continues daily to get the TAM and FSS attitude control solution 
operational.  The UARS testing and training simulator (UTTS) is being used for 
all testing.  In order to insure the validity of the test the FOT has been 
debugging some problems with the simulator attitude models.  


                            MEETINGS

The FOT attended the UARS Flight Software CCB on 20 November 2000.

The FOT attended a biweekly UARS Attitude meeting on 21 November 2000.


                     FUTURE SIGNIFICANT EVENTS

The FOT will put the TAM into monitor mode on 28 November 2000.  This will 
allow for the FOT to observe the TAM's performance and make any adjustments 
before using the TAM as an input to the update filter.

The next UARS Yaw slew is scheduled for December 02, 2000 (DOY 337).  This 
maneuver will be a Reverse to Forward Flight maneuver (Northern Hemisphere 
viewing).


Attachments: Appendix A,  Real-time Commands Sheets
             Appendix B,  Out-Of-Limits & ALERT Occurrences
             Appendix C,  Clock & Solar Array Rate Corrections
             Appendix D,  UARS Battery Performance Data
             Appendix E,  UARS Anomaly Reports

                            APPENDIX A
                    REAL-TIME COMMAND SHEETS
                  Week ending 24 November 2000

ORBIT     AOS        NO.       SUBSYS    REASON
-----     ---        ---       ------    ------
50232     323/1813   00-651    NBTR      NOOPS D-324
50237     324/0249   00-652    OBC       Clk Adj 20.6 to 0.00 msec/day
50239     324/0528   00-651    NBTR      NOOPS D-324
50247     324/1733   00-651    NBTR      NOOPS D-324
50247     324/1814   00-653    NBTR      NOOPS D-325
50250     324/2325   00-654    PEM       AXIS POWER OFF
50252     325/0212   00-655    HALOE     Power ON
50254     325/0530   00-653    NBTR      NOOPS D-325
50262     325/1734   00-653    NBTR      NOOPS D-325
50262     325/1814   00-656    NBTR      NOOPS D-326
50265     325/2329   00-657    OBC       Clk Adj 0.00 to 20.6 msec/day
50268     326/0435   00-656    NBTR      NOOPS D-326
50275     326/1554   00-656    NBTR      NOOPS D-326
50278     326/1922   00-659    NBTR      NOOPS D-327
50283     327/0437   00-659    NBTR      NOOPS D-327
50284     327/0534   00-659    NBTR      PLAYBACK
50287     327/1051   00-659    NBTR      PLAYBACK (FDF)
50288     327/1131   00-660    OBC       Clk Adj 20.6 to 0.00 msec/day
50289     327/1415   00-659    NBTR      NOOPS D-327
50292     327/1815   00-661    NBTR      NOOPS D-328
50298     328/0438   00-661    NBTR      NOOPS D-328
50304     328/1416   00-661    NBTR      NOOPS D-328
50307     328/1817   00-662    NBTR      NOOPS D-329
50317     329/1054   00-663    HALOE     CLEAR CONTROL STATUS WORD 4
50318     329/1237   00-664    HALOE     RELOAD BORESIGHT VALUES
50321     329/1559   00-662    NBTR      NOOPS D-329
50323     329/1932   00-665    NBTR      NOOPS D-330
___________________________________________________________________________

                           APPENDIX B
               OUT-OF-LIMITS & ALERTS OCCURRENCES
                   Week ending 24 November 2000

The following Out-Of-Limits occurrences were detected:

MNEMONIC     STATE     DESCRIPTION                    ORBITS
--------     -----     -----------                    ------
HACS4        Y-HI      HALOE CONTROL STATUS WORD 4    50317


The following ALERT occurrences were detected:

MNEMONIC     DESCRIPTION     ORBITS
--------     -----------     ------
              -None this report period

_______________________________________________________________________________
                            APPENDIX C
               CLOCK & SOLAR ARRAY RATE CORRECTIONS
                  Week ending 24 November 2000

                  CLOCK DRIFT RATE ADJUSTMENTS

ORBIT     TIME           CLOCK RATE CHANGE
-----     -----          -----------------
50237     324/025200     Clk Adj from 20.6 to 0.00 msec/day     
50265     325/233300     Clk Adj from 0.00 to 20.6 msec/day     
50288     327/113230     Clk Adj from 20.6 to 0.00 msec/day     

__________________________________________________________________________

                           APPENDIX D
               UARS BATTERY PERFORMANCE DATA
                    Week ending 24 Nov 2000


     17 Nov 2000 - GMT Day 322 Beta = 60.8 deg, SA Pos = 269 deg

     DIFF V (mV)    TEMP (C)     CURR (amp)   EON V  AVE C/D   AVE
BAT  MAX/MIN        MAX/MIN      MAX/MIN      MIN    RATIO     DOD(%)
---  -----------    --------     --------     -----  ------    ------
1   +728.0/+728.0   -3.05/-4.31  +0.0/+0.0    14.2   0.000     0.0
2   +11.2/-22.4     +3.87/+2.60  +23.2/-14.4  26.7   1.030     16.8
3   +11.2/-33.6     +2.60/+1.66  +24.4/-13.6  26.7   1.020     16.4


     24 Nov 2000 - GMT Day 329 Beta = 32.8 deg, SA Pos = 269 deg

     DIFF V (mV)    TEMP (C)     CURR (amp)   EON V  AVE C/D   AVE
BAT  MAX/MIN        MAX/MIN      MAX/MIN      MIN    RATIO     DOD(%)
---  -----------    ----------   --------     -----  ------    ------
1    +728.0/+728.0  -3.05/-4.62  +0.0/+0.0    14.2   0.000     0.0
2    +11.2/-56.0    +5.79/+4.51  +29.6/-15.2  25.1   1.030     21.1
3    +28.0/-33.6    +4.51/+3.23  +30.4/-16.0  25.0   1.030     20.7


     V/T LEVEL CHANGES
ORB#      TIME           LVL FR-TO
-----     ----           ----------
          -no changes    V/T 5
_________________________________________________________________________


                              APPENDIX E
                     UARS ANOMALY REPORTS
                  Week ending 24 November 2000

                   NEW ANOMALY REPORTS GENERATED
AR NO.     SUBSYS     ORBIT     TIME     TTR     PROBLEM/DATA LOSS
------     ------     -----     ----     ---     -----------------
00-054     FOT        50280   326/2147   23187   RESET PTP /2:39 32K


             ANOMALY CLOSURE INFORMATION RECEIVED
AR NO.     SUBSYS     ORBIT     TIME     TTR     PROBLEM/DATA LOSS
------     ------     -----     ----     ---     -----------------
            -None this report period


            ANOMALY INVESTIGATION REPORTS GENERATED
AR NO.     SUBSYS     ORBIT     TIME     TTR     PROBLEM/DATA LOSS
------     ------     -----     ----     ---     -----------------
           -None this report period

                     OPEN ANOMALY REPORTS
AR NO.     SUBSYS    ORBIT     TIME      TTR     PROBLEM/DATA LOSS
------     ------    -----     ----      ---     -----------------
00-047     FOT       see list  Nov 00    14170   Generic Late Acquisition
00-053     FOT       50168     319/1113  23157   Late Acquisition
00-054     FOT       50280     326/2147  23187   RESET PTP /2:39 32K


	 	SEVNTFW00/47